OpenStudy (anonymous):

Take the derivative of sin^2[lnx]

OpenStudy (anonymous):

\[f(x)=\sin^2(\ln(x))\] like that?

OpenStudy (anonymous):

or \[f(x)=\sin^2(x)\ln(x)\]

OpenStudy (anonymous):

first one requires the chain rule twice second one the chain and produce rule

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

ok first visualize it as \[\huge \left(\sin(\ln(x))\right)^2\]

OpenStudy (anonymous):

since that is what the square means on the sine then work from the outside in using the chain rule the derivative of something squared is two times something, times the derivative of the something

OpenStudy (anonymous):

the derivative of the sine of something is the cosine of something times the derivative of something

OpenStudy (anonymous):

and the derivative of the log is \(\frac{1}{x}\)

OpenStudy (anonymous):

I got this: 2(sin[lnx])(cos[lnx})(1/x)

OpenStudy (anonymous):

looks good to me

OpenStudy (anonymous):

Can I simplify that even more? @satellite73

OpenStudy (anonymous):

there is no such mathematical operation as "simplify" but no you cannot do anything with it unless you want to rewrite is as \[\frac{2\sin(\ln(x))\cos(\ln(x))}{x}\]

OpenStudy (anonymous):

oh okay. Thanks for your help! :)

OpenStudy (anonymous):

yw